The number of roots of the equation √(x-2) (x2 – 4x + 3) = 0 is
    (A)    one
    (B)    two
    (C)    three
    (D)    four

Grade:11

1 Answers

Sourabh Singh IIT Patna
9 years ago
Hii


Three roots (x-1) =0
(x-2) =0
(x-3) =0


thanks.


.
.
….
.
.
.
..
.
.
.
..
.

.
.